OpenStudy (anonymous):

Quotient of 4 + 6i/ 2 - 3i a.-5/26 - 6/13i b. - 10/13 + 24/13i c. 2 -2i d. None

OpenStudy (anonymous):

\[\frac{4+6i}{2-3i}\cdot\frac{2+3i}{2+3i}=\frac{8+12i+12i+18i^2}{4-9i^2}=\frac{8-18+24i}{4+9}=\frac{-10+24i}{13}\]

OpenStudy (anonymous):

so B?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you so much. I did this on my calculator and chose D. Your explanation helps.

OpenStudy (anonymous):

yw